# NICER Magnetar Burst Catalog

### v1

@@ 3.5 Correlation
4 Summary
In this study, we constructed a comprehensive catalog of magnetar short bursts spanning eight years of NICER observations. We identified a total of 1130 bursts, making this the largest burst catalog to date. Most of the bursts originated from SGR 1935+2154, which contributed 865 events. We present the distributions of timing and spectral parameters for all detected bursts and investigate their correlations. The bursts from SGR 1935+2154 generally show longer durations than those from other magnetars. We also report the first detection of s-scale bursts in magnetars, one from 1E 1048.15937 and another from CXOU J010043.1721134.
@@ Appendix A The Catalog
This appendix presents all bursts identified in this study. Magnetars for which at least one burst was detected include CXOU J010043.1721134 (Table 5), 4U 0142+61 (Table 6), 1E 1048.15937 (Table 7), Swift J1555.25402 (Table 8), 1RXS J170849.0400910 (Table 9), SGR 180620 (Table 10), XTE J1810197 (Table 11), Swift J1818.01607 (Table 12), Swift J1822.31606 (Table 13), SGR 18300645 (Table 14), 1E 1841045 (Table 15), SGR 1935+2154 (Table 16), 1E 2259+586 (Table 17), and the magnetar-like pulsar PSR J18460258 (Table 18). For magnetars with more than 30 detected bursts, a complete burst catalog is provided as an online table.

### v2

@@ III.5 Correlation
IV Summary
In this study, we constructed a comprehensive catalog of magnetar short bursts spanning eight years of NICER observations. We identified a total of 1130 bursts and most of these bursts originated from SGR 1935+2154, which contributed 865 events. We present the distributions of timing and spectral parameters for all detected bursts and investigate their correlations. The bursts from SGR 1935+2154 generally show longer durations than those from other magnetars. We also report the first detection of s-scale bursts in magnetars, one from 1E 1048.15937 and another from CXOU J010043.1721134.
@@ Appendix A The Catalog
This appendix presents all bursts identified in this study. Magnetars for which at least one burst was detected include CXOU J010043.1721134 (Table A1), 4U 0142+61 (Table A2), 1E 1048.15937 (Table A3), Swift J1555.25402 (Table A4), 1RXS J170849.0400910 (Table A5), SGR 180620 (Table A6), XTE J1810197 (Table A7), Swift J1818.01607 (Table A8), Swift J1822.31606 (Table A9), SGR 18300645 (Table A10), 1E 1841045 (Table A11), SGR 1935+2154 (Table A12), 1E 2259+586 (Table A13), and the magnetar-like pulsar PSR J18460258 (Table A14). For magnetars with more than 30 detected bursts, a complete burst catalog is provided as an online table.